OpenStudy (anonymous):

Anyone knows much about logarithmic equations? I need help. log base 3(x^2-4x-18)=3. Or if you can only help me to factor out the quadratic part would help too.

OpenStudy (anonymous):

Write your log equation in exponential form. So you get 3^3 = x^2 - 4x - 18 Or, 27 = x^2 -4x - 18. Set the quadratic equal to zero, and solve for x. Make sure you check that the value(s) of x work in the ORIGINAL equation.

OpenStudy (isaiah.feynman):